OpenStudy (anonymous):

WILL MEDAL (SAT) scores are normally distributed with a mean of 500 points and a standard deviation of 100 points. Suppose you take the SAT and several weeks later you receive a letter telling you that your results on the math portion of the test were in the 90th percentile. Recalling that SAT scores are always expressed as multiples of 10, how many points did you get on the test? A 690 B 600 C 590 D 130 E 630

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

sorry i don't know any statistics at all

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@karatechopper

OpenStudy (anonymous):

@Kainui

jimthompson5910 (jim_thompson5910):

do you have a calculator?

OpenStudy (anonymous):

No

jimthompson5910 (jim_thompson5910):

go here https://www.prometric.com/en-us/clients/SOA/Pages/calculator.aspx and type in 0.9 into the box next to N(x) under the "INVERSE CDF CALCULATOR"

jimthompson5910 (jim_thompson5910):

this will give you the value of k such that \(\Large P(Z \le k) = 0.9\) is true

OpenStudy (anonymous):

Is says x is 1.28155

jimthompson5910 (jim_thompson5910):

So that means \[\Large P(Z \le 1.28155) \approx 0.9\]

jimthompson5910 (jim_thompson5910):

we know that z = 1.28155

jimthompson5910 (jim_thompson5910):

mu = 500 is the population mean sigma = 100 is the population standard deviation

jimthompson5910 (jim_thompson5910):

z = 1.28155 mu = 500 sigma = 100 plug those values into z = (x - mu)/sigma and solve for x

OpenStudy (anonymous):

X=628.155

jimthompson5910 (jim_thompson5910):

round that to the nearest tens place

OpenStudy (anonymous):

630?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

you do that because "SAT scores are always expressed as multiples of 10"

OpenStudy (anonymous):

Oh ok thank you

jimthompson5910 (jim_thompson5910):

no problem